OpenStudy (shrutipande9):

a culture of bacteria grows at a rate which is inversely propertional to the square root of the number present. if there are 9 units present initially and 16 unit after 2 days. find in how many days the bacteria will be 36 units?

hartnn (hartnn):

so, \(\large a_n=\dfrac{k}{\sqrt{a_{n-1}}}\) with a1 = 9 got this equation ?

hartnn (hartnn):

with a2 = 16, n=2 you could get the value of k

OpenStudy (shrutipande9):

haann...yes ok

OpenStudy (shrutipande9):

ok cool...got it..thnx!:D

hartnn (hartnn):

welcome ^_^

OpenStudy (shrutipande9):

i got n as 10.21 days....is it correct?

hartnn (hartnn):

what u got k as ?

OpenStudy (shrutipande9):

fir se ainvyi medal huh

OpenStudy (shrutipande9):

i got k as 12.33

hartnn (hartnn):

a2 = 16 = k/ sqrt 9 didn't u use this ?

OpenStudy (shrutipande9):

see...i also need 2 find d constant c...

OpenStudy (shrutipande9):

see wat i was doing...i took dN/dt=k/sqr rt N here N is ur a

hartnn (hartnn):

ok

OpenStudy (shrutipande9):

so i need 2 integrate dis thing rite

hartnn (hartnn):

right, calculus stuff.... should have known

OpenStudy (shrutipande9):

so i m in d rite direction na?

hartnn (hartnn):

yes, go on

OpenStudy (shrutipande9):

so then after integrating and subs t=0 and N=9 i got c=18

hartnn (hartnn):

yup,

OpenStudy (shrutipande9):

ok cool..so in d same manner N=16 t=2 so k=12.33 and den i reached t=10.21days

hartnn (hartnn):

yes, i too get 10,219

OpenStudy (shrutipande9):

one more thing int of....sqr rt dN/dt=2(sqr rtN)^3/3 RITE?

hartnn (hartnn):

yes, because we divide by the power, which is here 3/2

OpenStudy (shrutipande9):

COOL.....thnx

hartnn (hartnn):

welcome ^_^

OpenStudy (shrutipande9):

btw aj ye mera saval tha..not my sister's:P

hartnn (hartnn):

cool! but m equally happy to help you both :)

OpenStudy (shrutipande9):

:) maths guru

hartnn (hartnn):

guru bhai guru bhai avya che..dhoom dhadaka karela che :P

OpenStudy (shrutipande9):

guru bhai ekdum maja ma che:)